If the utility function involving commodities X and Y is U= 8x + 22y - x² - 2xy - y² and the budjet constraint is 12= x + 2y, determine the values of x and y that maximize the utility?

1
Expert's answer
2022-02-01T06:55:22-0500

The values of x and y maximize the utility, if:

MUx/Px = MUy/Py or MUx/1 = MUy/2 and x + 2y = 12.

8 - 2x - 2y = 11 - x - y,

y = -x - 3,

x - 2x - 6 = 12,

x = -18,

y = 15.